# Visibility-Aware Keypoint Localization for 6DoF Object Pose Estimation

### v1

@@ 3.2 Handling Visibility of Symmetric Objects
To enforce the consistency of keypoint visibility labels, we can modify the annotated pose to maximize the number of internally visible keypoints in a fixed subset of keypoints $\mathcal{P}_{\mathrm{sym}}$. We consider internal visibility $V_{\mathrm{in}}$ because self-occlusion always exists and is not effected by external objects. For an object with discrete symmetry, before the whole training process, we obtain $\mathcal{P}_{\mathrm{sym}}$ by finding the largest visible subset under sampled poses. In practice, we uniformly sample 2,562 rotation matrices in $SO(3)$, and use a fixed translation $\mathbf{t}$ along the camera looking direction. Then during training, we can enumerate the finite equivalent poses to find the one maximizing internally visible keypoints in $\mathcal{P}_{\mathrm{sym}}$. For an object with continuous symmetry, we further derive an analytic solution. Without loss of generality, we assume $z$-axis is a symmetry axis. We modify the original annotated rotation matrix by right-multiplying rotation around $z$-axis $\mathbf{R}_{z}(\theta)$, where $\theta$ is the rotation angle around $z$-axis, and is determined by
@@
where $a,b$ are the first and second entry of $\mathbf{R}^{\top}\mathbf{t}$, respectively, []delimited-[][$\cdot$][ ⋅ ] is Iverson bracket outputting binary values $\{0,1\}$. We provide more details including the derivation in the Supplementary Material.
3.3 Visibility-Aware Importance via Personalized PageRank
So far we have discussed how to assign binary labels for keypoint visibility. While visible points are clearly more useful than invisible ones for pose estimation, not all visible points are equally important, depending on the distribution of these points and the specific estimation algorithms. Moreover, if an invisible keypoint is close to visible ones, localizing it may not be difficult. To address these issues, we propose to compute real-valued visibility-aware importance by measuring the closeness to visible keypoints w.r.t. a specific measure of the proximity.
In practice, we adopt Personalized PageRank (PPR) as our proximity measure, which is a variation of PageRank [40] based on a random walk model. Specifically, we build a directed $k$-nearest neighbor ($k$-NN) graph $G$ from predefined keypoints $\mathcal{P}$. For each keypoint $\mathbf{p}$, we create edges from $\mathbf{p}$ to its $k$-nearest neighbors. We define transition matrix $\mathbf{T}$ as
@@ 3.3 Visibility-Aware Importance via Personalized PageRank
where $N_{\mathrm{vis}}$ denotes the number of visible keypoints. By rearranging Eq. 8, we can compute $\mathbf{r}$ as
@@
Note that $\mathbf{I}-c\mathbf{T}$ is invertible thus $\mathbf{T}_{\mathrm{PPR}}$ is well-defined in Eq. 11. Moreover, $\mathbf{T}_{\mathrm{PPR}}$ is invariant for the object, so we can precompute $\mathbf{T}_{\mathrm{PPR}}$ using Eq. 11 and store it, and then we can obtain our visibility-aware importance $\mathbf{r}$ from any input binary labels using only one matrix-vector multiplication (Eq. 10).
Besides the efficient computation of importance $\mathbf{r}$, PPR also utilizes the overall structure of the object via the random walker on $G$. In contrast, other popular measures may focus on limited structure information, e.g., shortest paths. As a result, PPR is less impacted by outliers in visible keypoints, which makes the computation of $\mathbf{r}$ more robust.
3.4 From Visibility-Aware Importance to Pose Estimation
Figure 3: Illustration of our visibility-aware pose estimator. (a) Overall architecture: We use a backbone network to extract image features from the input image, and select keypoints with high visibility-aware importance for localization. We generate supervision signals online for keypoint visibility and localization. We also train object segmentation with available ground truth masks (omitted in the figure). (b) Details of visibility-aware selection module: For each keypoint, we use a multi-label classifier to predict external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$. We then obtain overall visibility $V$, and adopt Personalized PageRank (PPR) to generate visibility-aware importance from $V$. We finally sort importance and select keypoints with high importance.
To improve 6DoF pose estimation, we propose to localize keypoints with high visibility-aware importance $\mathbf{r}$ (Eq. 10). One feasible way is to train a network to directly regress $\mathbf{r}$. However, precisely producing importance of each keypoint is difficult. We instead obtain $\mathbf{r}$ in an alternative way. We first use binary classifiers to predict binary visibility $V$ (Sec. 3.1) and compute restart vector $\mathbf{s}$ (Eq. 9), then we get importance $\mathbf{r}$ using Eq. 10. In this way, we only need to train binary classifiers, which is typically easier than training regression models. Besides, as discussed in Sec. 3.3, our proposed PPR-based importance is robust against outliers in visibility estimation.
@@ 3.4 From Visibility-Aware Importance to Pose Estimation
Our proposed visibility-aware importance $\mathbf{r}$ can be easily integrated into existing keypoint-based 6DoF pose estimators. We adopt the recent state-of-the-art method CheckerPose [31] as our base. CheckerPose utilizes GNNs to explicitly model the interactions among dense keypoints, and predicts binary codes as a hierarchical representation of 2D locations. Besides, CheckerPose exploits a CNN decoder to learn image features and fuses the features into the GNN branch. To facilitate mini-batch-based training for GNNs, we select $N^{\prime}=N/2$ keypoints with highest importance for localization. For extreme case when the ratio of estimated visible keypoints is below a certain threshold, we directly use $N^{\prime}$ evenly distributed keypoints for robustness.
@@
As shown in Fig. 3, we use a backbone network to extract image features from input region of interest (RoI) $I_{O}$ of object $O$. Then we utilize a multi-label classifier to predict external visibility $V_{\mathrm{ex}}(\mathbf{p})$ and internal visibility $V_{\mathrm{in}}(\mathbf{p})$ for each keypoint $\mathbf{p}\in\mathcal{P}$. Each decomposed visibility focuses on a specific factor as discussed in Sec. 3.1. Within the classifier, we adopt a shallow GNN on $k$-NN graph $G$ described in Sec. 3.3, to improve the classification consistency for connected keypoints. After estimating $V_{\mathrm{ex}}(\mathbf{p}),V_{\mathrm{in}}(\mathbf{p})$, we can get overall binary visibility $V(\mathbf{p})$ by simply multiplying $V_{\mathrm{ex}}(\mathbf{p})$ and $V_{\mathrm{in}}(\mathbf{p})$. From the predicted binary visibility of each keypoint, we generate real-valued visibility-aware importance using PPR based algorithm (Eq. 10). The coefficient matrix $\mathbf{T}_{\mathrm{PPR}}$ is precomputed using Eq. 11, so we only need to create restart vector $\mathbf{s}$ from binary visibility (Eq. 9) and run a simple matrix-vector multiplication. We then select $N^{\prime}=N/2$ keypoints with highest importance, denoted as $\mathcal{P}_{\mathrm{vis}}\subset\mathcal{P}$. We use the subgraph $G_{\mathrm{vis}}$ of $G$ induced by $\mathcal{P}_{\mathrm{vis}}$ in GNN-based localization process.
@@
We further enhance keypoint embedding to improve the network performance. Specifically, We use a shallow GNN to obtain $64$-dim embedding $F_{P}$ from the coordinates of $\mathcal{P}$, and concatenate it with initial keypoint embedding in $G$. We can regard $F_{P}$ as positional encoding [37, 57], which facilitates the localization of the dynamic subgraph $G_{\mathrm{vis}}$.
3.5 Two-stage Training
We train our visibility-aware pose estimator in a supervised manner. As shown in Fig. 3, with object-level annotations, we generate binary labels of visibility described in Sec. 3.1 on the fly during training. Note that other supervision signals are either directly available (e.g., segmentation masks) or can also be efficiently generated on the fly (e.g., binary codes representing 2D locations).
@@ 3.5 Two-stage Training
While the whole network can be trained from scratch, it is often inefficient since the graph $G_{\mathrm{vis}}$ may vary significantly for the same input image $I$ when visibility prediction is unstable. We instead use a two-stage training procedure inspired by CheckerPose [31]. At the first stage, we train network layers corresponding to low-level estimations, including external visibility $V_{\mathrm{ex}}$, internal visibility $V_{\mathrm{in}}$, 1-bit indicator code $\mathbf{b_{v}}$, and the first $d_{0}$ bits of $\mathbf{b_{x}},\mathbf{b_{y}}$ ($\mathbf{b_{v}},\mathbf{b_{x}}$ and $\mathbf{b_{y}}$ are defined in CheckerPose). We treat the prediction of these low-level quantities as binary classification tasks, and use binary cross-entropy loss for training. At the second stage, we train the whole network. We use binary cross-entropy loss for visibility estimation and binary code generation, and apply $L_{1}$ loss for segmentation mask prediction.
4 Experiments
4.1 Experimental Setup
Implementation Details. We implement our method using PyTorch [43]. We train our network using the Adam optimizer [25] with a batch size of 32 and learning rate of 2e-4. We adopt CheckerPose [31] as our base keypoint localization method, and follow it to use $N=512$ predefined keypoints and $k=20$ nearest neighbors for $k$-NN graph. We select $N^{\prime}=256$ keypoints and use the induced subgraph in the localization step. For the binary code generation, we set the number of bits $d$ as 7, and the number of initial bits $d_{0}$ as 3. We use Progressive-X [2] to obtain pose from the correspondences for all experiments.
@@ 4.1 Experimental Setup
Evaluation Metrics. We adopt the commonly-used metric ADD(-S) to evaluate the estimated poses. To compute ADD(-S) with threshold $x\%$, we transform the 3D model points using the predicted poses and the ground truth, compute the average distance between the transformed results, and check whether the average distance is below $x\%$ of the object diameter. For symmetric objects, we compute the average distance based on the closest points. On YCB-V, we follow the common practice [64, 54, 33] to report the AUC (area under curve) of ADD-S and ADD(-S) [68], where the symmetric metric is used for all objects in ADD-S but for symmetric objects only in ADD(-S). The AUC metric reflects the accumulated performance across different thresholds, and provides complementary evaluation to ADD(-S). For keypoint-based methods, we compute average localization error by comparing the predicted keypoint locations and ground truth. We also report the inlier ratio with threshold of $x$ pixels.
4.2 Visibility Prediction
Table 1: Evaluation of visibility prediction. We report average accuracy, precision, recall, and F1 score of external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$ on three datasets.
Dataset $V_{\mathrm{ex}}$ $V_{\mathrm{in}}$
Acc. Prec. Recall F1 Acc. Prec. Recall F1
LM [11] 92.6 94.2 97.7 95.9 98.1 98.2 97.9 98.0
LM-O [3] 87.4 84.6 85.6 84.6 91.7 91.4 91.7 91.5
YCB-V [68] 91.8 94.2 94.1 94.0 95.8 95.6 95.7 95.6
We first demonstrate that visibility classification is a relatively easy task to enable reliable keypoint selection for localization. For the prediction of external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$, we compute accuracy, precision, recall, and F1 score. All metrics are evaluated on the networks after 50k training steps in the first stage. We report average results on the three datasets in Tab. 1. As shown in Tab. 1, a simple classifier is capable of producing good estimations.
4.3 Keypoint Localization
Table 2: Evaluation of keypoint localization. We compare our method with CheckerPose [31]. We report the average localization error, and inlier ratio with two thresholds (2 pixels and 5 pixels).
Dataset LM [11] LM-O [3] YCB-V [68]
[31] Ours [31] Ours [31] Ours
Error (pixel) 3.4 3.1 14.4 14.4 10.9 8.7
Inlier-2px (%) 44.8 67.4 19.4 24.3 8.4 15.2
Inlier-5px (%) 88.4 88.4 67.8 71.7 39.6 55.2
We evaluate keypoint localization in Tab. 2 and compare with our baseline method CheckerPose [31]. For the LM dataset where the occlusions are mild, we significantly increase the inlier ratio with the stricter threshold (2 pixels) and achieve comparable performance with threshold of 5 pixels. For the LM-O and YCB-V datasets with severe occlusions and clutters, we significantly boost the inlier ratio. The results clearly show that our visibility-aware strategy improves the quality of 3D-2D correspondences.
4.4 Ablation Study on LM Dataset
Table 3: Ablation Study on the LM Dataset.

### v5

@@ III-B Handling Visibility of Symmetric Objects
For an object with discrete symmetry, before the whole training process, we obtain $\mathcal{P}_{\mathrm{sym}}$ by finding the largest visible subset under sampled poses. In practice, we uniformly sample 2,562 rotation matrices in $SO(3)$, and use a fixed translation $\mathbf{t}$ along the camera looking direction. Then during training, we can enumerate the finite equivalent poses to find the one maximizing internally visible keypoints in $\mathcal{P}_{\mathrm{sym}}$.
For an object with continuous symmetry, we can further derive analytic formulas that are easy to evaluate. With Eq. (4) and Eq. (5), we get
$\begin{split}-\mathbf{d}_{\mathbf{p}}^{(c)}\cdot\mathbf{n}_{\mathbf{p}}^{(c)}&=(\mathbf{R}\mathbf{p}+\mathbf{t})^{\top}(\mathbf{R}\mathbf{n}_{\mathbf{p}})\\ &=\mathbf{p}^{\top}\mathbf{R}^{\top}\mathbf{R}\mathbf{n}_{\mathbf{p}}+\mathbf{t}^{\top}\mathbf{R}\mathbf{n}_{\mathbf{p}}\\ &=\mathbf{p}^{\top}\mathbf{n}_{\mathbf{p}}+(\mathbf{R}^{\top}\mathbf{t})^{\top}\mathbf{n}_{\mathbf{p}},\end{split}$ (6)
and we can see that the first term $\mathbf{p}^{\top}\mathbf{n}_{\mathbf{p}}$ is invariant. To maximize internally visible keypoints in $\mathcal{P}_{\mathrm{sym}}$, we only need to minimize the second term $(\mathbf{R}^{\top}\mathbf{t})^{\top}\mathbf{n}_{\mathbf{p}}$ for $\mathcal{P}_{\mathrm{sym}}$.
Base case: single continuous symmetry axis, no additional discrete symmetry. The symmetry transformations can be parameterized as $\mathbf{R}_{*}(\theta)$, where * denotes a symmetry axis and $\theta$ is the rotation angle around the symmetry axis. For a specific transformation $\mathbf{R}_{*}(\theta)$, the corresponding transformed rotation is
$\tilde{\mathbf{R}}=\mathbf{R}\mathbf{R}_{*}(\theta),$ (7)
and the corresponding transformed translation remains the same, thus
$(\tilde{\mathbf{R}}^{\top}\tilde{\mathbf{t}})^{\top}\mathbf{n}_{\mathbf{p}}=(\mathbf{R}^{\top}\mathbf{t})^{\top}(\mathbf{R}_{*}(\theta)\mathbf{n}_{\mathbf{p}}).$ (8)
Without loss of generality, we can assume $z$-axis is the symmetry axis. Otherwise, we can apply a simple coordinate system transformation to the object to satisfy this assumption. Consider a special case when $\mathbf{n}_{\mathbf{p}}=\mathbf{n}_{0}=[1,0,0]^{\top}$, and then
$\mathbf{R}_{z}(\theta)\mathbf{n}_{0}=\begin{bmatrix}\cos{\theta}&-\sin{\theta}&0\\ \sin{\theta}&\cos{\theta}&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}\cos{\theta}\\ \sin{\theta}\\ 0\end{bmatrix},$ (9)
and we further get
$(\tilde{\mathbf{R}}^{\top}\tilde{\mathbf{t}})^{\top}\mathbf{n}_{0}=a\cos{\theta}+b\sin{\theta},$ (10)
where $a,b$ are the first and second entry of $\mathbf{R}^{\top}\mathbf{t}$, respectively. For simplicity, denote $(\tilde{\mathbf{R}}^{\top}\tilde{\mathbf{t}})^{\top}\mathbf{n}_{0}$ as $f(\theta)$, and now we need to minimize $f(\theta)$. When $a=0$, it is straightforward to get $\theta=\pi/2+[b>0]\pi$, where []delimited-[][$\cdot$][ ⋅ ] is Iverson bracket outputting binary values $\{0,1\}$. When $a\neq 0$, we have
$f^{\prime}(\theta)=-a\sin\theta+b\cos\theta,$ (11)
and $\theta=\arctan{(b/a)}$ is one of the solutions to make $f^{\prime}(\theta)=0$. We also need to consider
$f^{\prime\prime}(\theta)=-a\cos\theta-b\sin\theta=-a\cos\theta(1+\frac{b^{2}}{a^{2}}),$ (12)
and let $f^{\prime\prime}(\theta)\geq 0$. Thus when
@@
$(\tilde{\mathbf{R}}^{\top}\tilde{\mathbf{t}})^{\top}\mathbf{n}_{0}$ achieves its minimum value, thus the quantity in Eq. (6) also achives its minimum value for the point $\mathbf{p}_{0}$ with $\mathbf{n}_{\mathbf{p}}=\mathbf{n}_{0}=[1,0,0]^{\top}$ and its neighborhood, which ensures the internal visibility of $\mathbf{p}_{0}$ and its neighborhood (i.e., $\mathcal{P}_{\mathrm{sym}}$).
In summary, we modify the annotated rotation matrix by right-multiplying rotation around $z$-axis $\mathbf{R}_{z}(\theta)$, where $\theta$ is the rotation angle and is determined by Eq. (13).
Variation I: single continuous symmetry axis, additional discrete symmetry. We can first Eq. (13) to resolve the ambiguity caused by continuous symmetry, then enumerate the finite equivalent poses of the discrete symmetry.
Variation II: multiple continuous symmetry axes. Without loss of generality, we can assume $y$-axis is another symmetry axis. After resolving the ambiguity caused by continuous symmetry of $z$-axis, we apply additional transformation $\mathbf{R}_{y}(\phi)$, where $\phi$ is the rotation angle around $y$-axis. The derivation of $\phi$ is very similar to $\theta$ in $\mathbf{R}_{z}(\theta)$, hence we omit it for simplicity.
Figure 2: Illustration of our visibility-aware pose estimator. (a) Overall architecture: We use a backbone network to extract features from the input image, and select keypoints with high visibility-aware importance for localization. We generate supervision signals online for keypoint visibility and localization. We also train object segmentation with available ground truth masks. We finally recover the pose using a PnP solver. (b) Details of visibility-aware selection module: For each keypoint, we use a multi-label classifier to predict external visibility $V_{\mathrm{ex}}$ and internal visibility $V_{\mathrm{in}}$. We then obtain overall visibility $V$, and adopt Personalized PageRank (PPR) to generate visibility-aware importance from $V$.
III-C Visibility-Aware Importance via Personalized PageRank
We can further compute real-valued visibility-aware importance by measuring the closeness to visible keypoints w.r.t. a specific measure of the proximity. In practice, we adopt Personalized PageRank (PPR) as our proximity measure, which is a variation of PageRank [23] based on a random walk model. Specifically, we build a directed $k$-nearest neighbor ($k$-NN) graph $G$ from predefined keypoints $\mathcal{P}$. For each keypoint $\mathbf{p}$, we create edges from $\mathbf{p}$ to its $k$-nearest neighbors. We define transition matrix $\mathbf{T}$ as
@@ III-C Visibility-Aware Importance via Personalized PageRank
where $N_{\mathrm{vis}}$ denotes the number of visible keypoints. By rearranging Eq. (15), we can compute $\mathbf{r}$ as
@@
Note that $\mathbf{T}_{\mathrm{ppr}}$ is well-defined in Eq. (18) because $\mathbf{I}-c\mathbf{T}$ is invertible. Moreover, $\mathbf{T}_{\mathrm{ppr}}$ is invariant for the object, so we can precompute $\mathbf{T}_{\mathrm{ppr}}$ using Eq. (18) and store it.
III-D From Visibility-Aware Importance to Pose Estimation
@@
We first use binary classifiers to predict binary visibility $V$ (Section III-A) and compute restart vector $\mathbf{s}$ (Eq. (16)), then we get importance $\mathbf{r}$ using Eq. (17). In this way, we only need to train binary classifiers, which is typically easier than training regression models of $\mathbf{r}$. Our proposed visibility-aware importance $\mathbf{r}$ can be easily integrated into existing keypoint-based 6DoF pose estimators. We adopt the recent state-of-the-art open-source method CheckerPose [16] as our base. CheckerPose utilizes GNNs of dense keypoints and predicts binary codes as a hierarchical representation of 2D locations. The 2D locations can be easily obtained from this representation without voting schemes. Besides, CheckerPose exploits a CNN decoder to learn image features and fuses the features into the GNN branch. To facilitate mini-batch-based training for GNNs, we select $N^{\prime}=N/2$ keypoints with highest importance for localization. For extreme case when the ratio of estimated visible keypoints is below a certain threshold, we directly use $N^{\prime}$ evenly distributed keypoints for robustness.
@@
As shown in Figure 2, we use a backbone network to extract image features from input region of interest (RoI) $I_{O}$ of object $O$. Then we utilize a multi-label classifier to predict external visibility $V_{\mathrm{ex}}(\mathbf{p})$ and internal visibility $V_{\mathrm{in}}(\mathbf{p})$ for each keypoint $\mathbf{p}\in\mathcal{P}$. Within the classifier, we adopt a shallow GNN on $k$-NN graph $G$ described in Section III-C. We can get overall binary visibility $V(\mathbf{p})$ by simply multiplying $V_{\mathrm{ex}}(\mathbf{p})$ and $V_{\mathrm{in}}(\mathbf{p})$, and generate real-valued visibility-aware importance $\mathbf{r}$ using PPR based algorithm (Eq. (17)). We then select $N^{\prime}=N/2$ keypoints with highest importance, denoted as $\mathcal{P}_{\mathrm{vis}}\subset\mathcal{P}$. We use the subgraph $G_{\mathrm{vis}}$ of $G$ induced by $\mathcal{P}_{\mathrm{vis}}$ in GNN-based localization process.
@@
We further enhance keypoint embedding to improve the network performance. Specifically, We use a shallow GNN to obtain $64$-dim embedding $F_{P}$ from the coordinates of $\mathcal{P}$, and concatenate it with initial keypoint embedding in $G$. We can regard $F_{P}$ as positional encoding [38], which facilitates the localization of the dynamic subgraph $G_{\mathrm{vis}}$.
@@
We use a two-stage training procedure to train our visibility-aware pose estimator. At the first stage, we train network layers corresponding to low-level estimations, including external visibility $V_{\mathrm{ex}}$, internal visibility $V_{\mathrm{in}}$, 1-bit indicator code $\mathbf{b_{v}}$, and the first $d_{0}$ bits of $\mathbf{b_{x}},\mathbf{b_{y}}$ ($\mathbf{b_{v}},\mathbf{b_{x}}$ and $\mathbf{b_{y}}$ are defined in CheckerPose). At the second stage, we train the whole network. We use binary cross-entropy loss for visibility estimation and binary code generation, and apply $L_{1}$ loss for segmentation mask prediction.
III-E Extension for CAD-Free Object Pose Estimation
When the 3D CAD model is unavailable, we can use an off-the-shelf image-based mesh reconstructor. The biggest problem is the imperfect reconstruction quality, especially the inaccurate reconstruction of invisible regions. To address this issue, for each object, we use $m$ views to reconstruct $m$ 3D meshes. For each reconstructed mesh, we use the criteria in Section III-A to determine the visible vertices in the corresponding input view, from which we uniformly sample $N/m$ keypoints. In this way, we obtain predefined $N$ keypoints with plausible 3D coordinates and normals. Furthermore, when generating synthetic training images, for each randomly sampled pose, we render the mesh reconstructed from the closest view. Once we generate sufficient training images, we can train VAPO following the CAD-based setting. When prior knowledge of symmetry is unavailable, we assume that all objects are asymmetric during training.
IV Experiments
IV-A Experimental Setup
@@
Implementation Details. We implement our method using PyTorch [41]. We train our network using the Adam optimizer [42] with a batch size of 32. Following [16], we train a unified pose estimator for all 13 LM objects, while training separate pose estimators for each object on LM-O and YCB-V. The training follows a two-stage procedure. In the first stage, we train the layers generating low-level quantities for 50k steps with a learning rate of 2e-4. In the second stage, for LM, we continue training all network layers for 100k steps, followed by an additional 20k steps with a reduced learning rate of 1e-4. We train all network layers for 700k steps on LM-O and 250k steps on YCB-V, both using a learning rate of 2e-4. We adopt CheckerPose [16] as our base keypoint localization method, and follow it to use $N=512$ predefined keypoints and $k=20$ nearest neighbors for $k$-NN graph. We select $N^{\prime}=256$ keypoints and use the induced subgraph in the localization step. For the binary code generation, we set the number of bits $d$ as 7, and the number of initial bits $d_{0}$ as 3. We apply data augmentation and off-the-shelf object detectors following [18, 16]. To obtain final poses, we run Progressive-X for 400 iterations with a reprojection error threshold of 2 pixels. For the CAD-free setting, we use Wonder3D [43] to reconstruct $m=8$ meshes from $m=8$ views.
Evaluation Metrics. We adopt the commonly-used metric ADD(-S) to evaluate the estimated poses. To compute ADD(-S) with threshold $x\%$, we transform the 3D model points using the predicted poses and the ground truth, compute the average distance between the transformed results, and check whether the average distance is below $x\%$ of the object diameter. For symmetric objects, we compute the average distance based on the closest points. On YCB-V, we also report the AUC (area under curve) of ADD-S and ADD(-S) [26], where the symmetric metric is used for all objects in ADD-S but for symmetric objects only in ADD(-S).
IV-B Ablation Study on LM Dataset
TABLE I: Ablation Study on LM. We highlight the best result and the second best result in red and blue, respectively.